OpenStudy (daisy.xoxo619):

Lines perp? line 1 w/ coordinates (0,1) (6,-2) line 2 w/ coordinates (3,-3) (6,3) What I Know so far : slope line 1= -1/2 slope line 2=2

OpenStudy (daisy.xoxo619):

@mathstudent55

OpenStudy (mathstudent55):

Good, you are correct so far. Now multiply the two slopes. If the answer is -1, the lines are perpendicular. That's what it means for the slopes to be "negative reciprocals", it's that their slopes have a product of -1.

OpenStudy (daisy.xoxo619):

ok how would i multiply the two slopes?

OpenStudy (mathstudent55):

Change the second slope into a fraction by setting it over 1. Then multiply the fractions by multiplying the numerators together and the denominators together. \(-\dfrac{1}{2} \times 2 = -\dfrac{1}{2} \times \dfrac{2}{1} = -\dfrac{1 \times 2}{2 \times 1} = -\dfrac{2}{2} = -2\)

OpenStudy (daisy.xoxo619):

so would the solution be there perp, bcuz product of there slopes =-1

OpenStudy (mathstudent55):

Yes. That is exactly it.